OpenStudy (anonymous):

The Sum of first N consecutive odd integers is N^2. What is the sum of all odd integers between 13 and 39 inclusive. A. 351 B. 364 C. 410 D. 424 E. 450

OpenStudy (anonymous):

\[39^2-11^2\]

OpenStudy (anonymous):

i.e. add from 1 to 39, and subtract from1 to 11

OpenStudy (shubhamsrg):

it wont be 39^2 - 11^2 ,,by N we mean the nth odd term.. so out ans should be 20^2 - 6^2

OpenStudy (shubhamsrg):

@satellite73

OpenStudy (raden):

yeah, it will becomes 20^2-6^2 = (20+6)(20-6) = 26 * 14 :)